OpenStudy (anonymous):

Marianne is debating between two different mortgages for $175,000. She found a 20-year fixed rate loan at 7.55% and 15-year fixed rate loan at the same rate. How much more interest will she pay for the 20-year loan versus the 15-year loan?

OpenStudy (anonymous):

p=175000 rate r=7.55% time t=20-15=5years more interest paid=prt/100=175000*7.55*5/100=66062.500